C = 250-0.75Y
I = 350
G = 800
X = 500
M = 200+0.4Y
T = 0.15Y
Use the above information to answer the following question
If C = 250-0.75Y, interpret the meaning of the value 250
State the value of MPT, MPC and MPS on disposable income and MPM
Find MPC and MPS on national income
Find the level of income in the economy
Calculate the net exports and interpret
If investment rise by 15%, what will be the change in national income
1
Expert's answer
2019-10-22T09:36:56-0400

C is autonomous consumption, i.e. consumption that is independent of income.


Y = C + I + G + (X − M)  = 250 - 0.75Y+350+800+(500-200-0.4Y) = 1400 -0.75Y+500-200-0.4Y = 1700 - 1.1Y

2.1Y=1700

Y=1700/2.1=809.5


net exports = Х-М = 500-200-0.4Y=300-0.4*809.5=-2938. this means that imports exceed exports.


MPC = С/Y = (250-0.75Y)/809.5 = 0.44

MPS = 1-0.44 = 0.56


if investments increase by 15%, total income will increase by more than 15% due to the effect of the investment multiplier